A luxury ship can accommodate 1,000 passengers together with its crew. It has 20 life boat in case of an emergency. each life boat (raft) measures 3 m x 2.5 m x 60 cm. if each raft must accommodate 70 % only of its capacity, what should be the maximum height or depth of the raft be submerged into the sea water. How many 50 kg people can each life boat accommodate? can the 20 rafts be sufficient for the passengers and crew? if not, what should be the suggestion you can give? (assume all other relevant data were considered in the design of the life boats. density of sea water is 1.03 x3 kg/cubic m.)


1
Expert's answer
2021-02-18T18:44:00-0500

"a = 3m \space b = 2.5m \space c = 0.6m"

"V = abc= 3*2.5*0.6=4.5m^3"


"F_A = Nm_og"

"\rho_sV_sg = Nm_og \to N = \large\frac{\rho_sV_s}{m_o}" "=\large\frac{\rho_s0.7V_s}{m_o}" "=\large\frac{\rho_s0.7abc}{m_o}" "= \large\frac{1030*0.7*3*2.5*0.6}{60}" "= 64.89"

at least 64 men

that is to say each boat can save 64 men

20 * 64 = 1320 is enough for 1000 passengers